OpenStudy (anonymous):

A flea jumps by exerting a force of 1.27*10^-5 N straight down on the ground. A breeze blowing on the flea parallel to the ground exerts a force of 1.20*10^-6 N on the flea. Find the direction and magnitude of the acceleration of the flea if its mass is 6.0*10^-7 kg. (Let us assume that Fwind points to the right. We will consider this to be the +x direction and vertical to be the +y direction.)

OpenStudy (aaronq):

So there are 3 forces acting on the flea  Write newton's second law for these \(\sum F_x=F_{wind}=ma_x \rightarrow a_x=\dfrac{F_{wind}}{m}\) \(\sum F_y=F_{jump}+F_g=F_{jump}-mg=ma_y \rightarrow a_y=\dfrac{F_{jump}}{m}-g\) So the acceleration vector are: \(a_x=\dfrac{F_{wind}}{m}\) \(a_y=\dfrac{F_{jump}}{m}-g\) Now it's just trigonometry you need